# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

### v2

@@ Abstract
Index Terms:
Distributed model predictive control, privacy preservation, differential privacy.
@@ I Introduction
Over the past decades, model predictive control (MPC) has enjoyed great success due to its ability to explicitly handle system constraints and guarantee prescribed control performance [1, 2]. MPC can be implemented either in a centralized or distributed manner. Centralized MPC relies on a central unit to process all system information and solve the online optimization problem, which often results in poor scalability and requires substantial computation power, especially for complex or large-scale systems. Consequently, distributed MPC (DMPC) has garnered much attention in recent years, offering many advantages of distributed systems and proving to be an effective tool for various applications, including vehicle platoons [3], microgrids [4], and multi-robot systems [5].
Based on the type of couplings between subsystems, DMPC studies can be roughly divided into three categories, i.e., coupling in cost functions, coupling in system dynamics, and coupling in constraints. In this paper, we focus on systems with coupled global constraints, which have many real-world applications [6, 7]. Several approaches have been proposed to guarantee the strict satisfaction of coupled constraints in a distributed manner. In [8], a sequential DMPC method is developed, which first divides the global problem into several local subproblems and then solves them sequentially in a given order. The satisfaction of coupled constraints is guaranteed through plan exchanges among subsystems. [9] proposes a DMPC approach for flexible communication that can remove optimization order restrictions. Building on [8] and [9], [10] presents a parallel DMPC approach allowing simultaneous optimizations and maintaining flexible communication and parallel computation advantages. However, global optimality remains unclear in these approaches. In [11], a dual decomposition-based DMPC approach is designed to explicitly pursue global optimality. This approach transforms the dual problem of DMPC into a consensus optimization problem and then solves it using the distributed alternating direction multiplier method (ADMM). To improve the convergence speed when solving DMPC, a Nesterov-accelerated-gradient algorithm is utilized in [12]. In addition, a push-sum dual gradient algorithm and a noisy ADMM algorithm are developed in [13] and [14] to address the DMPC problem under time-varying directed communication network and communication noise, respectively.
In the aforementioned methods [11, 12, 13, 14], distributed optimization algorithms are leveraged to address the DMPC problem, requiring each subsystem/agent to share explicit local information with neighboring subsystems/agents in order to satisfy coupled global constraints. Note that the shared messages often contain sensitive information, which raises significant concerns about privacy leakage. For instance, an eavesdropper could wiretap the communication channel and deduce privacy-sensitive information from exchanged messages. When DMPC is adopted in specific domains like smart grid or intelligent transportation, the disclosure of privacy-sensitive information may further pose safety risks and lead to economic losses. Considering the growing awareness of privacy and security, it is imperative to ensure privacy protection in DMPC. So far, few results are available on the privacy preservation of DMPC, but privacy-preserving approaches for distributed optimization have been well studied. For the latter, one typical technique is (partially) homomorphic encryption, which has been utilized in existing works such as [15, 16]. The encryption-based approaches use cryptography to conceal privacy-sensitive information and can be directly extended for the privacy protection of DMPC [17]. However, this technique generally requires tedious encryption and decryption procedures, leading to huge overheads in both communication and computation. Another technique relies on spatially or temporally correlated noises/uncertainties [18, 19, 20], aiming to obscure information shared in distributed optimization. Due to the correlated nature of these noises/uncertainties, these approaches are typically vulnerable to adversaries with access to all messages shared in the communication network.
With implementation simplicity and rigorous mathematical foundations, differential privacy (DP) has witnessed growing popularity, emerging as a de facto standard for privacy protection. In recent years, DP-based privacy methods have been introduced in distributed optimization by integrating DP noise into objective functions [21] or exchanged information [22, 23, 24]. Nevertheless, the direct injection of persistent DP noise to existing algorithms inevitably compromises optimization performance, resulting in an inherent trade-off between accuracy and privacy. It is crucial to note that extending DP-based privacy approaches to DMPC is not straightforward, as the compromise on optimization accuracy can deteriorate control performance and potentially lead to constraint violations.
@@
In this paper, a differentially-private DMPC algorithm is designed for linear discrete-time systems with coupled global constraints. We first demonstrate the need for privacy preservation by showcasing that a conventional distributed dual gradient algorithm for DMPC is vulnerable to eavesdropping attacks. A DP noise injection mechanism is then introduced into the distributed dual gradient algorithm, which obscures the private information exchanged among subsystems to prevent adversaries from inferring sensitive information. By leveraging the results in [25, 26], a weakening factor sequence and a step-size sequence are carefully designed to effectively mitigate the influence of DP noise. Rigorous analysis shows that the proposed algorithm can ensure almost sure convergence to a global optimal solution and maintain $\epsilon$-differential privacy with a finite cumulative privacy budget. Aligned with the privacy-preserving distributed algorithm, we provide an implementation strategy for DMPC, ensuring the recursive feasibility and stability of the closed-loop system. Simulations are performed to validate the performance of the proposed scheme.
@@
The rest of this paper is organized as follows. In Section II, the preliminaries of DMPC and differential privacy are introduced. In Section III, a new differentially-private distributed dual gradient algorithm is developed, and convergence analysis is conducted. Section IV presents the implementation strategy of DMPC. Finally, a numerical study is given in Section V, and concluding remarks are summarized in Section VI.
@@
Notations: $\mathbb{R}^{n}$ stands for the $n$-dimensional Euclidean space. Given two integers $a$ and $b$ ($a<b$), $\mathbb{Z}_{a}^{b}$ represents the set $\{a,a+1,\cdots,b\}$. $I_{n}$ denotes the identify matrix of dimension $n$. ${\bf 1}_{n}$ and ${\bf 0}_{n}$ represent the $n$-dimensional column vector with all entries being 1 and 0, respectively. We use $Q>(\geq)0$ to denote that $Q$ is a positive definite (semi-definite) matrix. $\|x\|$ and $\|x\|_{1}$ represent the standard Euclidean norm and the $L_{1}$ norm of a vector $x$, respectively. Moreover, $\|x\|_{Q}^{2}:=x^{\top}Qx$.
@@ II-A Problem Description
Consider $M$ linear discrete-time subsystems where each is described as follows:
$x_{i}(t+1)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\quad i\in\mathbb{Z}_{1}^{M}.$ (1)
In (1), $x_{i}(t)\in\mathbb{R}^{n_{i}}$ and $u_{i}(t)\in\mathbb{R}^{m_{i}}$ are the state and control input of subsystem $i$ at time instant $t$, respectively. The state and control input of subsystem $i$ should satisfy the following local constraints:
$x_{i}(t)\in\mathcal{X}_{i},\quad u_{i}(t)\in\mathcal{U}_{i},$ (2)
where $\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ and $\mathcal{U}_{i}\subset\mathbb{R}^{m_{i}}$ denote state and control input constraint sets, respectively. Moreover, all the subsystems are subject to $p$ global constraints described by
@@
where $\varPsi_{x_{i}}\in\mathbb{R}^{p\times n_{i}}$ and $\varPsi_{u_{i}}\in\mathbb{R}^{p\times m_{i}}$ are some given matrices.
@@
Each linear discrete-time subsystem, i.e., $(A_{i},B_{i})$, is controllable. Additionally, $\mathcal{X}_{i}$ and $\mathcal{U}_{i}$ are bounded and closed polytopes which contain the origins as their inner point.
In this paper, we consider the same DMPC problem as presented in [11, 12, 13, 14, 17]. Specifically, based on (1)-(3), the DMPC problem is formulated as
@@
$\;\mathrm{s.t.}\quad\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$ (4b)
$\quad\quad\sum_{i=1}^{M}f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})\leq b(\varepsilon).$ (4c)
@@
$J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})\!:=\!\sum_{\ell=0}^{N-1}\!\left(\|\tilde{x}_{i}(\ell|t)\|_{Q_{i}}^{2}\!+\!\|\tilde{u}_{i}(\ell|t)\|_{R_{i}}^{2}\right)\!+\!\|\tilde{x}_{i}(N|t)\|_{P_{i}}^{2},$ (5)
where $N\in\mathbb{Z}_{>0}$ is the length of prediction horizon, $\tilde{x}_{i}(\ell|t)$ and $\tilde{u}_{i}(\ell|t)$ are the $\ell$th step predicted state and control input at time instant $t$, respectively, $\tilde{{\bm{u}}}_{i}:=\{\tilde{u}_{i}(0|t),\cdots,\tilde{u}_{i}(N-1|t)\}$ stands for the predicted input sequence over the prediction horizon, and $Q_{i}>0$, $R_{i}>0$, and $P_{i}>0$ are weight matrices. For each subsystem $i$, $P_{i}$ is the solution of the following algebraic Riccati equation:
@@
$\tilde{\mathcal{U}_{i}}(x_{i}(t)):=\{\tilde{{\bm{u}}}_{i}\in\mathbb{R}^{m_{i}N}:$ (7)
$\tilde{x}_{i}(\ell+1|t)=A_{i}\tilde{x}_{i}(\ell|t)+B_{i}\tilde{u}_{i}(\ell|t),\tilde{x}_{i}(0|t)=x_{i}(t),$
$\tilde{x}_{i}(\ell|t)\in\mathcal{X}_{i},\tilde{u}_{i}(\ell|t)\in\mathcal{U}_{i},\tilde{x}_{i}(N|t)\in\mathcal{X}_{i}^{f},\ell\in\mathbb{Z}_{0}^{N-1}\},$
with $\mathcal{X}_{i}^{f}$ being the terminal constraint set. In addition, the global coupled constraint in (4c) is a tightened form of the constraint in (3), and $f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ and $b(\varepsilon)$ are given by
$f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ $:=\begin{bmatrix}\varPsi_{x_{i}}\tilde{x}_{i}(0|t)+\varPsi_{u_{i}}\tilde{u}_{i}(0|t)\\ \vdots\\ \varPsi_{x_{i}}\tilde{x}_{i}(N-1|t)+\varPsi_{u_{i}}\tilde{u}_{i}(N-1|t)\end{bmatrix},$ (8)
$b(\varepsilon)$ $:=\begin{bmatrix}(1-\varepsilon M){\bf 1}_{p}\\ \vdots\\ (1-\varepsilon MN){\bf 1}_{p}\end{bmatrix},$
where $0\leq\varepsilon<\frac{1}{MN}$ is a tolerance parameter. The introduction of the tightened constraint is to ensure that the numerical algorithm used to solve the DMPC problem can be terminated in advance. To facilitate the feasibility and stability analysis of DMPC, the terminal constraint set $\mathcal{X}_{i}^{f}$ can be chosen as a closed maximal polytope such that for any $x_{i}\in\mathcal{X}_{i}^{f}$, we have
$K_{i}x_{i}\in\mathcal{U}_{i},\quad(A_{i}+B_{i}K_{i})x_{i}\in\mathcal{X}_{i}^{f},$ (9)
$\sum_{i=1}^{M}(\varPsi_{x_{i}}+\varPsi_{u_{i}}K_{i})x_{i}\leq(1-\varepsilon MN){\bf 1}_{p}.$
For more details about the tightening of constraint (3) and the construction of the terminal constraint set $\mathcal{X}_{i}^{f}$, please refer to [11].
@@
For the initial system state $\{x_{1}(0),\cdots,x_{M}(0)\}$, the Slater condition holds, i.e., there exists $\{\tilde{{\bm{u}}}_{1},\cdots,\tilde{{\bm{u}}}_{M}\}$ that satisfies (4b) and (4c).
The communication network of $M$ subsystems is described by an interaction weight matrix $L=\{L_{ij}\}\in\mathbb{R}^{M\times M}$. Specifically, for each subsystem $i$, the neighbor set $\mathcal{N}_{i}$ consists of all subsystems $j$ that can directly communicate with subsystem $i$. If $j\in\mathcal{N}_{i}$, then $L_{ij}>0$; otherwise, $L_{ij}=0$. We define $L_{ii}:=-\sum_{j\in\mathcal{N}_{i}}L_{ij}$ for all $i\in\mathbb{Z}_{1}^{M}$. Moreover, $L$ satisfies the following assumption:
@@
The interaction weight matrix $L=\{L_{ij}\}$ is symmetric and satisfies ${\bf 1}_{M}^{\top}L={\bf 0}_{M}^{\top}$, $L{\bf 1}_{M}={\bf 0}_{M}$, and $\|I_{M}+L-\frac{{\bf 1}_{M}{\bf 1}_{M}^{\top}}{M}\|<1$.

### v3

@@ I Introduction
Over the past decades, model predictive control (MPC) has achieved great success due to its ability to explicitly handle system constraints and ensure desired control performance [Mayne2014AUTO]. MPC can be implemented in either a centralized or distributed manner. Centralized MPC requires a central unit to process all system information, making it computationally intensive and less scalable for large-scale systems. Consequently, distributed MPC (DMPC) has emerged as a promising alternative, offering the advantages of distributed systems and having been effectively applied in various areas [Hans2018TSE, Luis2020RAL].
DMPC studies can be roughly categorized based on the type of couplings between subsystems: cost function couplings, system dynamics couplings, and constraint couplings. This paper focuses on systems with coupled global constraints, which have many real-world applications [BaiTVT2023, AlizadehSPM2012, Notarstefano2019FT]. Existing methods address coupled constraints through techniques like sequential optimization [Richards2007IJC] and parallel computation [Trodden2014SCL]. However, global optimality remains unclear in these approaches. To achieve global optimality, a DMPC scheme based on distributed alternating direction multiplier method (ADMM) is proposed in [Wang2017AUTO]. Other extensions include a push-sum dual gradient algorithm for time-varying directed networks [Jin2020TAC], a noisy ADMM algorithm for handling communication noise [Li2021AUTO], and a primal-dual algorithm designed with contraction theory [SuTCYB2022].
The aforementioned methods [Wang2017AUTO, Jin2020TAC, Li2021AUTO, SuTCYB2022] employ distributed optimization to address DMPC problems, requiring subsystems to share local information to meet coupled global constraints. However, these shared messages and global constraints often contain sensitive data, raising concerns about privacy leakage. Eavesdroppers could wiretap communications and deduce private information, potentially leading to safety risks and economic losses. For example, in DMPC for automated vehicles [BaiTVT2023], global constraints are formulated with each vehicle’s position and velocity, which are sensitive and should be protected from disclosure. Similarly, in demand-side management for smart grids [AlizadehSPM2012], global constraints capture the relationship between aggregated customer loads and the power bid, potentially revealing proprietary consumption patterns. As such, ensuring privacy protection in DMPC is essential for both security and practical deployment. While few results address privacy in DMPC, privacy-preserving methods for distributed optimization are well established. For the latter, a common technique is homomorphic encryption, which conceals sensitive information via cryptography [Lu2018AUTO, Zhang2018TCNS] and can be extended to DMPC [Zhao2023TCNS]. However, this technique generally incurs high communication and computation overhead due to complex encryption and decryption processes. In contrast, differential privacy (DP) offers a lightweight alternative with strong theoretical guarantees. DP-based methods have been applied to distributed optimization by adding persistent noise to objective functions [Nozari2016TCNS] or shared information [Huang2015, Xiong2020TCNS, Ding2021TAC]. Nevertheless, the direct injection of DP noise to existing algorithms inevitably compromises optimization performance, resulting in a trade-off between accuracy and privacy. Note that extending DP-based methods to DMPC is particularly challenging, as the compromise on optimization accuracy can impair control performance and potentially violate constraints.
@@
In this paper, a differentially-private DMPC algorithm is designed for linear discrete-time systems with coupled global constraints. We first demonstrate the need for privacy preservation by showcasing that a conventional distributed dual gradient algorithm for DMPC is vulnerable to eavesdropping attacks. A DP noise injection mechanism is then introduced into the distributed dual gradient algorithm to obscure exchanged private information. Leveraging results from [WangTAC2024, WangTAC2023], we carefully design weakening factor and step-size sequences to effectively mitigate the influence of DP noise. Rigorous analysis shows that the proposed algorithm can ensure almost sure convergence to a global optimal solution and maintain $\epsilon$-differential privacy with a finite cumulative privacy budget. Aligned with the privacy-preserving distributed algorithm, we provide an implementation strategy for DMPC, ensuring the recursive feasibility and stability of the closed-loop system. Simulations are performed to validate the efficacy of the proposed scheme. Our work differs significantly from existing DMPC methods [Wang2017AUTO, Jin2020TAC, Li2021AUTO, SuTCYB2022] and DP-based privacy methods [Huang2015, Xiong2020TCNS, Ding2021TAC, WangTAC2024, WangTAC2023]. First, unlike [Wang2017AUTO, Jin2020TAC, Li2021AUTO, SuTCYB2022], which primarily focus on control design and are susceptible to eavesdropping, we integrate a novel noise injection mechanism into the DMPC framework to address both control and privacy issues. Second, in contrast to algorithms in [Huang2015, Xiong2020TCNS, Ding2021TAC] that directly add DP noise to exchanged information and have to trade convergence accuracy for privacy, our algorithm uses weakening factor and step-size sequences to minimize the adverse impact of DP noise, achieving both a finite cumulative privacy budget and guaranteed convergence. Third, [WangTAC2024, WangTAC2023] are designed for distributed optimization, and cannot be applied to DMPC problems. Our work borrows ideas from [WangTAC2024, WangTAC2023] to construct tailored weakening factor and step-size sequences, and develops a DMPC-specific implementation strategy to extend the applicability of these techniques to the distributed control framework.
@@
The rest of this paper is organized as follows. In Section II, the preliminaries of DMPC and DP are introduced. In Section III, a new differentially-private distributed dual gradient algorithm is developed, and convergence analysis is conducted. Section IV presents the implementation strategy of DMPC. Finally, a numerical study is given in Section V, and concluding remarks are summarized in Section VI.
@@
Notations: $\mathbb{R}^{n}$ stands for the $n$-dimensional Euclidean space, and $\mathbb{R}_{+}^{n}$ is the non-negative orthant of $\mathbb{R}^{n}$. Given two integers $a$ and $b$ ($a<b$), $\mathbb{Z}_{a}^{b}$ represents the set $\{a,a+1,\cdots,b\}$. $I_{n}$ denotes the identity matrix of dimension $n$. ${\bf 1}_{n}$ and ${\bf 0}_{n}$ represent the $n$-dimensional column vector with all entries being 1 and 0, respectively. We use $Q>(\geq)0$ to denote that $Q$ is a positive definite (semi-definite) matrix. $\|x\|$ and $\|x\|_{1}$ represent the standard Euclidean norm and the $L_{1}$ norm of a vector $x$, respectively. Moreover, $\|x\|_{Q}^{2}:=x^{\top}Qx$. The Euclidean projection of vector $x$ on a convex set $\mathcal{X}\subset\mathbb{R}^{n}$ is denoted by $\Pi_{\mathcal{X}}\left[x\right]:=\mathop{\rm argmin}_{h\in\mathcal{X}}\|h-x\|$.
@@ II-A Problem Description
Consider $M$ linear discrete-time subsystems, each described as follows:
$x_{i}(t+1)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\qquad i\in\mathbb{Z}_{1}^{M},$ (1)
where $x_{i}(t)\in\mathbb{R}^{n_{i}}$ and $u_{i}(t)\in\mathbb{R}^{m_{i}}$ are the state and control input of subsystem $i$ at time instant $t$, respectively. Each subsystem $i$ should satisfy local constraints $x_{i}(t)\in\mathcal{X}_{i}$ and $u_{i}(t)\in\mathcal{U}_{i}$, with $\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ and $\mathcal{U}_{i}\subset\mathbb{R}^{m_{i}}$ being state and input constraint sets, respectively. Moreover, all the subsystems are subject to $p$ global constraints described by
@@
where $\varPsi_{x_{i}}\!\in\!\mathbb{R}^{p\times n_{i}}$ and $\varPsi_{u_{i}}\!\in\!\mathbb{R}^{p\times m_{i}}$ are some given matrices. The coupled linear constraints arise in many practical multi-agent systems, such as safety limits in automated vehicles [BaiTVT2023] and demand constraints in smart grids [AlizadehSPM2012].
@@
Each subsystem, i.e., $(A_{i},B_{i})$, is controllable. Additionally, $\mathcal{X}_{i}$ and $\mathcal{U}_{i}$ are bounded and closed polytopes which contain the origins as their inner point.
The DMPC problem is formulated as [Wang2017AUTO, Jin2020TAC, Li2021AUTO, SuTCYB2022]
@@
$\;\mathrm{s.t.}\quad\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t)),\quad\sum_{i=1}^{M}f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})\leq b(\varepsilon).$ (3b)
@@
$J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})\!=\!\sum_{\ell=0}^{N-1}\!\left(\|\tilde{x}_{i}(\ell|t)\|_{Q_{i}}^{2}\!+\!\|\tilde{u}_{i}(\ell|t)\|_{R_{i}}^{2}\right)\!+\!\|\tilde{x}_{i}(N|t)\|_{P_{i}}^{2},$ (4)
[…]
@@
$\tilde{\mathcal{U}_{i}}(x_{i}(t))=\{\tilde{{\bm{u}}}_{i}\in\mathbb{R}^{m_{i}N}:\tilde{x}_{i}(\ell+1|t)=A_{i}\tilde{x}_{i}(\ell|t)+B_{i}\tilde{u}_{i}(\ell|t),$ (6)
$\tilde{x}_{i}(0|t)\!=\!x_{i}(t),\tilde{x}_{i}(\ell|t)\!\in\!\mathcal{X}_{i},\tilde{u}_{i}(\ell|t)\!\in\!\mathcal{U}_{i},\tilde{x}_{i}(N|t)\!\in\!\mathcal{X}_{i}^{f},\ell\!\in\!\mathbb{Z}_{0}^{N-1}\},$
with $\mathcal{X}_{i}^{f}$ being the terminal constraint set. In addition, the global coupled constraint in (3b) is a tightened form of the constraint in (2), and $f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ and $b(\varepsilon)$ are given by
$f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ $=\begin{bmatrix}\varPsi_{x_{i}}\tilde{x}_{i}(0|t)+\varPsi_{u_{i}}\tilde{u}_{i}(0|t)\\ \vdots\\ \varPsi_{x_{i}}\tilde{x}_{i}(N-1|t)+\varPsi_{u_{i}}\tilde{u}_{i}(N-1|t)\end{bmatrix}$, (7)
$b(\varepsilon)$ $=\begin{bmatrix}(1-\varepsilon M){\bf 1}_{p}^{\top},\cdots,(1-\varepsilon MN){\bf 1}_{p}^{\top}\end{bmatrix}^{\top}$,
where $0\leq\varepsilon<\frac{1}{MN}$ is a tolerance parameter. To facilitate the feasibility and stability analysis of DMPC, the terminal constraint set $\mathcal{X}_{i}^{f}$ is selected to satisfy
$K_{i}x_{i}\in\mathcal{U}_{i},\qquad(A_{i}+B_{i}K_{i})x_{i}\in\mathcal{X}_{i}^{f},$ (8)
$\sum_{i=1}^{M}(\varPsi_{x_{i}}\!+\!\varPsi_{u_{i}}K_{i})x_{i}\!\leq\!(1-\varepsilon MN){\bf 1}_{p},\,\forall x_{i}\in\mathcal{X}_{i}^{f},\forall i\in\mathbb{Z}_{1}^{M}.$
For further details on the constraint tightening in (7) and the construction of the terminal constraint set $\mathcal{X}_{i}^{f}$ to satisfy (8), please refer to [Wang2017AUTO].
@@
For the initial state $\{x_{1}(0),\cdots,x_{M}(0)\}$, the Slater condition holds, i.e., there exists $\{\tilde{{\bm{u}}}_{1},\cdots,\tilde{{\bm{u}}}_{M}\}$ that satisfies (3b).
The communication network of $M$ subsystems is described by an interaction weight matrix $L=\{L_{ij}\}\in\mathbb{R}^{M\times M}$. Specifically, for each subsystem $i$, the neighbor set $\mathcal{N}_{i}$ consists of all subsystems $j$ that can directly communicate with subsystem $i$. If $j\in\mathcal{N}_{i}$, then $L_{ij}>0$; otherwise, $L_{ij}=0$. We define $L_{ii}:=-\sum_{j\in\mathcal{N}_{i}}L_{ij}$ for all $i\in\mathbb{Z}_{1}^{M}$.
@@
The matrix $L$ is symmetric and satisfies ${\bf 1}_{M}^{\top}L={\bf 0}_{M}^{\top}$, $L{\bf 1}_{M}={\bf 0}_{M}$, and $\|I_{M}+L-\frac{{\bf 1}_{M}{\bf 1}_{M}^{\top}}{M}\|<1$.